OpenStudy (anonymous):

What is the line of symmetry for the graph of y = -3x2 + 12x - 11?

OpenStudy (campbell_st):

for a parabola \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2 \times a}\] in your question a = -3 and b = 12 just substitute and calculate the answer.